# QPanda3: A High-Performance Software-Hardware Collaborative Framework for Large-Scale Quantum-Classical Computing Integration

### v1

@@ IV-B Intermediate Representation for Quantum Program Transmission - OriginBIS
V Optimized Compilation of Quantum Circuits
@@ V-A Transpiler of QPanda3
V-B Starting with an Example
@@
from pyqpanda3.core import QProg, draw_qprog, RZ, X1, CZ, SWAP, set_print_options
from pyqpanda3.transpilation import Transpiler, generate_topology
@@
prog_1 = QProg()
prog_1 << RZ(3,1.57) << RZ(4,3.14) << RZ(5,6.28) << CZ(3,4) << CZ(4,5) << CZ(3,5)
set_print_options(2)
[…]
The output is
@@
topo = [[0,1],[1,2]]
@@
transpiler = Transpiler()
prog_level_0 = transpiler.transpile(prog_1, topo, {}, 0)
@@
V-C Qubit Mapping and Routing
@@ V-D1 Sabre
V-D2 Independent Optimization Steps
@@
V-D3 Pre-placement Optimization and Post-routing Optimization
@@
VI Device-Based Quantum Program Analysis
@@
VI-A Runtime State Information and Quantum Program Analysis
@@
VI-B Quantum Circuit Analysis for Device Performance
VI-B1 Overview
@@
from pyqpanda3.profiling import draw_circuit_features
from pyqpanda3.core import QCircuit, H, CNOT, X
circuit = QCircuit(4)
circuit << H(0) << CNOT(0, 1) << CNOT(1, 2) << X(3)
draw_circuit_features(circuit, True)
@@
VI-C Performance Analysis of Quantum Programs Based on Flow and Device Information
@@ VI-C1 An Example
from pyqpanda3.profiling import draw_circuit_profile
from pyqpanda3.core import QCircuit, H, CNOT, X
cir1 = QCircuit(4)
cir1 << H(0) << CNOT(0, 1) << CNOT(1, 2) << X(3) << X(3)
cir2 = QCircuit(4)
cir2 << X(1) << cir1
draw_circuit_profile(cir2, {’CNOT’: 200, ’H’: 60, ’X’: 30}, True)
@@
VI-C2 Comparison with Classical Program Analysis
@@
VI-C3 Output Analysis Results

### v2

@@ A-B Intermediate Representation for Quantum Program Transmission - OriginBIS
Appendix B Optimized Compilation of Quantum Circuits
@@ B-A Transpiler of QPanda3
B-B Starting with an Example
@@
1from pyqpanda3.core import QProg, draw_qprog, RZ, X1, CZ, SWAP, set_print_options
2from pyqpanda3.transpilation import Transpiler, generate_topology
Listing 1: An example Python program that evaluates the properties of quantum states.
@@
1prog_1 = QProg()
2prog_1 << RZ(3,1.57) << RZ(4,3.14) << RZ(5,6.28) << CZ(3,4) << CZ(4,5) << CZ(3,5)
3set_print_options(2)
Listing 2: An example Python program that evaluates the properties of quantum states.
The output is show as Fig 21
@@
1topo = [[0,1],[1,2]]
Listing 3: An example Python program that evaluates the properties of quantum states.
@@
1transpiler = Transpiler()
2prog_level_0 = transpiler.transpile(prog_1, topo, {}, 0)
@@
Listing 4: An example Python program that evaluates the properties of quantum states.
B-C Qubit Mapping and Routing
@@ B-D1 Sabre
B-D2 Independent Optimization Steps
@@
B-D3 Pre-placement Optimization and Post-routing Optimization
@@
Appendix C Device-Based Quantum Program Analysis
@@
C-A Runtime State Information and Quantum Program Analysis
@@
C-B Quantum Circuit Analysis for Device Performance
C-B1 Overview
@@
1from pyqpanda3.profiling import draw_circuit_features
2from pyqpanda3.core import QCircuit, H, CNOT, X
3circuit = QCircuit(4)
4circuit << H(0) << CNOT(0, 1) << CNOT(1, 2) << X(3)
5draw_circuit_features(circuit, True)
Listing 5: An example Python program that evaluates the properties of quantum states.
@@
C-C Performance Analysis of Quantum Programs Based on Flow and Device Information
@@ C-C1 An Example
1from pyqpanda3.profiling import draw_circuit_profile
2from pyqpanda3.core import QCircuit, H, CNOT, X
3cir1 = QCircuit(4)
4cir1 << H(0) << CNOT(0, 1) << CNOT(1, 2) << X(3) << X(3)
5cir2 = QCircuit(4)
6cir2 << X(1) << cir1
7draw_circuit_profile(cir2, {’CNOT’: 200, ’H’: 60, ’X’: 30}, True)
Listing 6: An example Python program that evaluates the properties of quantum states.
@@
C-C2 Comparison with Classical Program Analysis
@@
C-C3 Output Analysis Results